Item 6 Item 6 A bond with a coupon rate of 7.6% (assume it is paid once annually), maturing in 10 years at a value of $1,000 and a current market price of $820 will have a current yield of (Round your answer to 2 decimal places.) Multiple Choice 9.27% 10.27% 8.27% 9.77%
Answer #1

Information provided:

Par value= future value= $1,000

Coupon rate= 7.6%

Coupon payment= 0.076*1,000= $76

Time= 10 years

Current value= present value= $820

Current yield is calculated using the below formula:

Current Yield= Annual interest/Current price

                          = $76/ 820*100

                          = 0.0927*100

                          = 9.27%.

Hence, the answer is option a.

In case of any query, kindly comment on the solution.